how do i find all double roots of
s^3-27s=0 showing all the steps?

Answer by RAY100(1637)

s^3 -27s =0
.
s(s^2 -27) = 0
.
s=0
.
s^2 -27= 0
.
s^2 =27
.
s = +/- sqrt 27
.
s= +/- sqrt(9*3)
.
s= +/- sqrt 9 * sqrt3
.
s = +/- 3sqrt3
.
.
.
check,,s=0,,,,,0=0,,,ok
.
s=+3sqrt3...(3sqrt3)^3 -27(3sqrt3) =0,,,ok
.
s=-3sqrt3....(-3sqrt3)^3 -27(-3sqrt3) = 0,,,,,ok
.